OpenStudy (anonymous):

Find parametric equations for the tangent line to the curve with the given parametric equations at the specified point. : Point P(9, 0, 2) x=1+8√(t) y=t^3 - t z= t^3 + t

OpenStudy (p0sitr0n):

\[T(x(t),y(t),z(t))=\frac{r'(t)}{|r(t)|}\] Now, you need to find t. this involves solving 9=1+8 root(t) 0=t^3-t 2=t^3+t then compute T, and your line is now T(t)